OpenStudy (anonymous):

the side of a triangle are in the ratio 3 : 4 : 5. What is the length of each side if the perimeter of a triangle is 90 cm? A.) 22.5 cm, 30 cm, and 37.5 cm B.) 19.3 cm, 25.7 cm, and 32.1 cm C.) 7.5 cm, 11.5 cm, and 32.1 cm D.) 10.5 cm, 11.5 cm, and 12.5 cm

OpenStudy (***[isuru]***):

If the ratio is 3:4:5 then the sides must be like this as a fraction of perimeter 3/12 , 4/12 , 5/12 If A is the first side A =90 * (3/12) =22.5 if B is the second side B = 90*(4/12) = 30 If C is the third side C = 90*(5/12) = 37.5 so these r the lengths of sides.... hope this will help ya

OpenStudy (anonymous):

So the answer would be D ?

OpenStudy (***[isuru]***):

No the answer is "A". Didn't u get the calculations I did ?

OpenStudy (***[isuru]***):

Also there is an very easy method to do this question..... Add up the 3 lengths given in each choice and the total value of the 3 length should be equal to 90.... And u will see that only the values given in "A" will add up to form 90